### v2

# EgoDex: Learning Dexterous Manipulation from Large-Scale Egocentric Video

Abstract
Imitation learning for manipulation has a well-known data scarcity problem. Unlike natural language and 2D computer vision, there is no Internet-scale corpus of data for dexterous manipulation. One appealing option is egocentric human video, a passively scalable data source. However, existing large-scale datasets such as Ego4D do not have native hand pose annotations and do not focus on object manipulation. To this end, we use Apple Vision Pro to collect EgoDex: the largest and most diverse dataset of dexterous human manipulation to date. EgoDex has 829 hours of egocentric video with paired 3D hand and finger tracking data collected at the time of recording, where multiple calibrated cameras and on-device SLAM can be used to precisely track the pose of every joint of each hand. The dataset covers a wide range of diverse manipulation behaviors with everyday household objects in 194 different tabletop tasks ranging from tying shoelaces to folding laundry. Furthermore, we train and systematically evaluate imitation learning policies for hand trajectory prediction on the dataset, introducing metrics and benchmarks for measuring progress in this increasingly important area. By releasing this large-scale dataset, we hope to push the frontier of robotics, computer vision, and foundation models. EgoDex is publicly available for download at https://github.com/apple/ml-egodex.
1 Introduction
The “bitter lesson” [36] of recent breakthroughs in large language models and large vision models is that the simple recipe of supervised learning with vast amounts of data is far more effective than competing approaches. Two key challenges have prevented the application of the bitter lesson to the longstanding challenge of autonomous robot manipulation: (1) it is unclear what data should be collected, and (2) it is unclear how such data can be collected at the requisite scale.
The leading approach to data collection for robot imitation learning is teleoperation, in which human operators provide demonstrations by directly controlling robot hardware. Recent works such as Open X-Embodiment [26] and DROID [16] pioneer community-wide efforts to pool together hundreds of hours of robot teleoperation data. While such datasets can be used effectively for pretraining robot control policies, teleoperation is bottlenecked by physical robot operation, and it is unclear how to continue scaling this paradigm beyond its current size. Other works explore learning visual representations from existing in-the-wild Internet videos and images [32, 20]. In this case, while large-scale data is available, unstructured video data lacks the precise annotation necessary to learn dexterous manipulation.
We explore a middle path between the two: egocentric human video with paired 3D hand pose annotations. As suggested by recent work [15, 30], such an approach is passively scalable, similar to text and images on the Internet. Effectively learning from such data is critical in a future where wearable headsets and smart glasses may be omnipresent. Data is a crucial component for doing so; before AlexNet [17] must come ImageNet [34].
To this end, we introduce EgoDex: a large-scale dataset and benchmark for learning dexterous manipulation from large-scale egocentric video. EgoDex consists of 829 hours of 30 FPS video and paired skeletal data with a total of 90 million frames and 338000 task demonstrations across 194 tabletop manipulation tasks. To our knowledge, the EgoDex dataset is the largest and most diverse dataset of dexterous human manipulation to date.
There are several key properties of the proposed data that make it more suitable for dexterous manipulation than existing alternatives:
  • EgoDex is passively scalable, unlike robot teleoperation and other approaches that require deliberate effort for data collection. EgoDex suggests the human hand as a common embodiment, unlike teleoperation and other approaches that collect data that is only compatible with specific robot hardware platforms.
  • EgoDex has 30 FPS 1080p egocentric video with a wide field of view, capturing much of what a human sees while manipulating objects. It has precise and highly detailed 3D pose information for the user’s head, arms, wrists, and each joint of each finger from on-device SLAM and calibrated cameras, containing critical dexterous manipulation data unlike in-the-wild Internet videos and Ego4D [11].
  • EgoDex consists of extremely diverse behaviors beyond simple pick-and-place such as unscrewing a bottle cap, flipping pages of a book, and plugging a charger into a socket. It consists entirely of active manipulation, unlike existing large egocentric video datasets such as Ego4D.
We systematically evaluate imitation learning policies for hand trajectory prediction to assess the state of the art and identify challenges for future research. We hope that EgoDex will not only accelerate progress in robot manipulation but also be useful more broadly in applications such as augmented reality, computer vision, assistive prosthetics, and human-computer interaction.
2 Related Work
Dataset # Traj. # Tasks # Frames Lang. Annot. Cam. Ext. Dexterous Annot. Collection Method
RoboTurk [21] 2k 3 12M teleoperation
RoboNet [9] 162k n/a 15M scripted
BridgeData V2 [38] 60k 13 2M teleop+scripted
DROID [16] 76k 86 19M teleoperation
EgoMimic [15] 2k 3 0.4M egocentric video
EPIC-KITCHENS [8] 40k 125 12M egocentric video
HOI4D [19] 4k 54 2M egocentric video
Ego4D (HOI) [11] 89k n/a 21M egocentric video
EgoDex (ours) 338k 194 90M egocentric video
Table 1: Comparison of different robot manipulation datasets (above middle line) and human manipulation datasets (below middle line). Ego4D (HOI) considers the subset of Ego4D that involves hand-object interaction. EgoDex has the largest amount of trajectories, tasks, and frames by a large margin and has language annotation, camera extrinsics, and dexterous annotation. “Dexterous annotation” is defined here as labels for multi-finger hand poses, which does not include lower fidelity pose data like parallel jaw robot grippers or wrist-only tracking.
2.1 Large-Scale Manipulation Datasets
Several prior works introduce large-scale open-source robot teleoperation datasets including RoboTurk [21], BridgeData [38], RT-X [26], and DROID [16]. While such datasets contain up to hundreds of hours of valuable manipulation data, it is not clear how to scale the paradigm further than its present scale. Robot teleoperation is extremely labor-intensive and resource-constrained, requiring an operational physical robot and a human teleoperator actively controlling the robot to perform each desired task. Furthermore, it is not clear to what degree such datasets can generalize beyond the set of hardware embodiments and camera viewpoints with which they were collected, even when the datasets consist of samples collected across multiple different embodiments.
Other large-scale datasets such as Ego4D [11] and EPIC-KITCHENS [8] consist of egocentric video recording humans performing various activities. While such datasets are more scalable and not limited to particular hardware platforms, they typically do not focus on manipulation and do not have paired 3D annotations for dexterous manipulation.
There is also a large body of work that considers hand-object interaction [19, 2, 4]. While these datasets often do have 3D hand pose annotations, they are orders of magnitude smaller than EgoDex due to manual annotation processes. Moreover, their emphasis is primarily on grasping rather than diverse and long-horizon manipulation tasks.
2.2 Scalable Methods for Robot Data Collection
Recent work identifies the data scarcity problem in robot imitation learning and proposes innovative techniques for scalable data collection. Chi et al. [7] propose the “universal manipulation interface”: handheld grippers that enable human teachers to provide demonstrations without physical robots. Wang et al. [39] introduce a portable data collection system with motion capture gloves. Others propose collecting robot-free demonstrations by simulating robot hardware in augmented reality [6, 27, 23].
These approaches all face a similar pitfall: they require active data collection. While they may make it easier to collect data than teleoperation, human demonstrators must still be incentivized to intentionally collect the data. Such approaches face a significant uphill battle in approaching the scale of Internet datasets, where text and images are not deliberately collected but rather a passive byproduct of human interaction with the Internet.
2.3 Learning from Human Video
Video data is abundant on the Internet. Prior work explores representation learning on unstructured large-scale image and video data for pretraining visual encoders [32, 20] and grasp affordances [1] for downstream manipulation. However, raw unstructured video data faces a prohibitively large gap between its image distribution and that of a dexterous manipulation task. Moreover, such videos are not labeled with corresponding motor actions with which to train a policy.
One option is to postprocess the unstructured video data with 3D hand prediction networks such as HaMeR [28], recently explored by Ren et al. [33]. However, the prediction quality of these networks can suffer without multiple viewpoints and detailed knowledge of the camera extrinsics at all times, usually unavailable with raw Internet video. In contrast, the EgoDex dataset includes 3D head and hand tracking at the time of collection, where multiple cameras on the Vision Pro, known intrinsics and extrinsics, and a production-grade hand prediction network all contribute to precise annotation.
Most similar to our work is EgoMimic [15], which proposes the collection of egocentric video and paired 3D hand tracking. The primary difference is scale: while EgoMimic collects around 4 hours of data, we collect 829 hours with a much broader data and task distribution. We also collect more dexterous annotations, critical for downstream manipulation: 3D positions and orientations for the upper body including the head, shoulders, arms, and 25 joints in each hand, whereas EgoMimic collects only the wrist positions. Lastly, a preview of EgoDex appeared in Humanoid Policy $\sim$ Human Policy [30], which used a 3% subset of EgoDex for successful human-to-robot transfer.
3 EgoDex Dataset
Figure 1: Distribution of EgoDex dataset. Top: Distribution of distinct verbs, sorted by frequency. The horizontal axis are verbs of EgoDex. The orange plot is taken from DROID [16]. While many verbs in DROID are below the $10^{1}$ mark, most verbs in EgoDex are above the $10^{3}$ mark. Bottom: Distribution of distinct objects. The clustering is suggested by GPT-4.
3.1 Overview
The EgoDex dataset contains 829 hours of 1080p, 30 Hz egocentric video with 338000 episodes across 194 tasks. This is a total of 90 million frames (i.e., data samples). The full dataset takes 2.0 TB of storage on disk. We compare EgoDex to existing manipulation datasets in Table 1. EgoDex has almost an order of magnitude more data than the next largest dataset as well as higher task diversity. It also has language annotations, camera extrinsics, and dexterous annotations (Section 3.3). The dataset is publicly available under a CC-by-NC-ND license; see Appendix A.4 for access instructions.
3.2 Data Collection
All data is collected with Apple Vision Pro running visionOS 2. The high-resolution and high-frequency passthrough and wide field of view enable intuitive egocentric data collection, where the collector can observe the environment unobstructed as if with their own eyes, and the camera data records precisely what the collector sees without any pose offsets (unlike, for instance, a head-mounted camera). Production-grade pose tracking software enables natural demonstration with bare hands and without any additional hardware apparatus.
[…]
Figure 2: Left: Joints captured by EgoDex. Right: Examples of dexterous manipulation behaviors. Tracked fingertips are highlighted in distinct colors and show 0.5 seconds of motion before the current frame. From left to right, top to bottom, the tasks are: unzipping a ziploc bag, removing a book from a bookshelf, removing a screw from a fixture, folding a t-shirt, decluttering, opening a case, unscrewing a bottle cap, tying shoelaces, and washing a cup.
3.3 Modalities
The data consists of the following: 1) Egocentric RGB video with $1920\times 1080$ resolution at 30 Hz frequency. 2) Camera intrinsics and extrinsics at 30 Hz. 3) 3D position and orientation of all upper body joints (including 25 joints for each hand) at 30 Hz. 4) Confidence values for pose predictions at 30 Hz. 5) Natural language annotation of the manipulation.
The metadata annotated by data collectors includes the task name, a brief task description in natural language, details about the environment, and details about the object(s) that are manipulated. Since the metadata can be noisy, these fields are provided as input to GPT-4 [25], which combines this information into a single detailed natural language description.
Confidence values are scalars between 0 and 1 indicating the ARKit prediction confidence per skeletal joint. A confidence of zero indicates that the joint is fully occluded from view. See Appendix A.2 for a comprehensive list of all the joints and more information.
3.4 Task Types
EgoDex consists of 3 types of tasks:
  • Reversible tasks are pairs of tasks that are the inverse of each other. The distribution of final states for one task is within the distribution of initial states for its inverse. For example, connecting a charger to a device and removing a charger from the device.
  • Reset-free tasks are tasks with a final state distribution that falls within its own initial state distribution. For example, throwing a ball in the air and catching it (where gravity acts as the reset).
  • Reset tasks are tasks in which the environment must be reset to the initial state distribution after each demonstration.
Reversible and reset-free tasks enable a higher yield from data collection as they eliminate costly resets, which are not included in the recorded data.
3.5 Diversity
Prior works [16, 26] identify several potential axes of demonstration diversity: viewpoint diversity, task diversity, scene diversity, object diversity, and more. In EgoDex, the emphasis is diversity in dexterous manipulation behaviors. Tasks and objects vary such that the required dexterity ranges far beyond pick-and-place, the primary behavior in most robot teleoperation datasets. For example, tasks include tightening a screw, tying shoelaces, dealing cards, flipping pages, catching tennis balls, and slotting batteries. The task distribution covers a wide range of everyday household manipulation tasks that can be performed on a tabletop surface. There is also a significant amount of basic pick-and-place with diverse objects as well as the benchmark tasks from the FurnitureBench assembly benchmark [13]. The full list of 194 tasks is provided in Appendix  A.1.
To get a sense of the spread of the task distribution, as in prior work we plot the distribution of de-duplicated verbs in Figure 1. We observe that the distribution is much wider than prior works such as DROID [16], where a large fraction of verbs have less than $10^{1}$ demonstrations and sometimes only a single demonstration; in contrast, most of the verbs in EgoDex have more than $10^{3}$ demonstrations.
Still, the verb distribution does not capture the full diversity of manipulation behaviors or tasks. For example, “assemble” can involve radically different behaviors in the context of different objects and tasks. See Figure 2 for examples of different dexterous manipulation behaviors captured in the dataset.
While the scene diversity in EgoDex is limited to tabletop environments, the Cartesian product of scene and behavior is not the focus of our work, which focuses on behavioral diversity. Scene diversity can be introduced with modern visual data augmentation methods such as image-to-image generative models [42, 5].
4 EgoDex Benchmarks
4.1 Action Representation
Given the full set of skeletal joints in the EgoDex dataset, many action representations are possible: wrist positions, wrist orientation, positions of fingertips, and so on. Since we focus on dexterous manipulation, we choose a representation that captures sufficient bimanual dexterity. Specifically, the action $\mathbf{a_{t}}$ at time $t$ is represented as the 3D position of each wrist, the 6D orientation of each wrist (as suggested by [43]), and the 3D position of each fingertip. Thus, each action has a total dimensionality of 2 hands $\times$ (3 + 6 + (3 $\times$ 5 fingertips)) = 48. In practice, actions are predicted in chunks over a fixed time horizon. Poses are expressed in the current camera frame [15], and each action chunk is a relative trajectory [7].
4.2 Benchmarks
We propose two benchmark tasks for EgoDex. The first is dexterous trajectory prediction: from the egocentric image observations, skeletal joint poses, and natural language description, the task is to predict the trajectories of the hands for a given time horizon following the observations. Specifically, we seek to train the following estimator:
$f_{\theta}(\mathbf{o}_{0..t},\mathbf{s}_{0..t},l)=\mathbf{\hat{a}}_{t:t+H}$
where $\mathbf{o}_{0..t}$ are the egocentric image observations up to and including time $t$, $\mathbf{s}_{0..t}$ are skeletal pose observations up to and including time $t$, $l$ is a natural language description, $\mathbf{\hat{a}}_{t:t+H}$ is the predicted action chunk, and $H$ is the prediction horizon.
Since multimodality can be very severe in natural human motion, the second benchmark is inverse dynamics: from the image observations and skeletal poses up to time $t$ as well as a goal image observation at the end of the time horizon, the task is to predict the trajectories of the hands in between the start and end observations. In this case, we train the following estimator, which can be interpreted as a visually goal-conditioned policy:
$f_{\theta}(\mathbf{o}_{0..t},\mathbf{s}_{0..t},\mathbf{o}_{t+H},l)=\mathbf{\hat{a}}_{t:t+H}$
Each of these benchmarks are parameterized by prediction horizon $H$. For example, a short-horizon trajectory prediction task may set $H=30$ (1 second), while a more difficult long-horizon task may set $H=90$ (3 seconds).
Unlike typical robot hardware experiments that can vary across physical environments, the EgoDex benchmarks are fully reproducible with a fixed training and test set. We set aside 1% of the EgoDex dataset as a fixed held-out test set for evaluations, where the remaining 99% can be split across training and validation as desired.
4.3 Evaluation Metrics
Since trajectory prediction for natural human motion is inherently multimodal, evaluating a single predicted trajectory against the ground truth sample may be insufficient for measuring correctness. For example, for the simple task of placing a fruit in a basket, it could be placed at variable locations within the basket, moved at variable speeds, and moved in different but equally valid trajectories from the initial position to the basket.
Thus, for each benchmark task we evaluate performance with a “best of KK” metric. For each data point in the full test set, we sample the trained model $K$ times to capture different possible modes. We then compute the distance between the ground truth trajectory and the trajectory closest to it out of the $K$ samples, where “distance” is calculated as the Euclidean distance between predicted 3D keypoint positions and their ground truth 3D counterparts, averaged over each timestep in the predicted chunk and each of the 12 keypoints (i.e., the wrist and fingertips of each hand). Intuitively, this value can be interpreted as the average positional error in 3D space between ground truth and prediction in meters. The final value is averaged over the full test set. For deterministic models, the value is the same regardless of $K$; for stochastic models, the value improves as $K$ increases, as the model gets more chances to sample the ground truth mode.
5 Experiments
Figure 3: Model prediction visualizations for Dec + BC on test set images with a 2 second horizon. Blue trajectories are ground truth and red trajectories are predictions, where darker colors and closer to the current frame and lighter colors are further in the future. The points shown are the wrist and fingertip positions projected into the camera frame (a total of 12 trajectories).
We train and evaluate state-of-the-art imitation learning policies from the X-IL framework [14] on the benchmarks from Section 4. Specifically, we train two Transformer model architectures (encoder-decoder and decoder-only) and three policy representations (behavior cloning, denoising diffusion, and flow matching). We also run experiments to evaluate the effect of prediction horizon, visual goal-conditioning, dataset size, and model size. In total we train and evaluate 14 different models. We train all models for 50,000 gradient steps with a batch size of 2048 parallelized across 8 NVIDIA A100 GPUs. Additional training and model details are provided in Appendix A.3. The results are presented in Tables 2, 3, 4 and Figure 4 and summarized below.
Encoder-decoder architectures outperform decoder-only. In Table 2 we observe that all encoder-decoder (“EncDec”) models consistently outperform their decoder-only (“Dec”) counterparts by a small margin.
Avg Distance (m) Final Distance (m)
Model $K=1$ $K=5$ $K=10$ $K=1$ $K=5$ $K=10$
Dec + BC $0.045$ $0.045$ $0.045$ $0.062$ $0.062$ $0.062$
Dec + DDPM $0.053$ $0.044$ $0.041$ $0.071$ $0.050$ $0.044$
Dec + FM $0.052$ $0.042$ $0.040$ $0.071$ $0.049$ $0.043$
EncDec + BC $\mathbf{0.044}$ $0.044$ $0.044$ $\mathbf{0.060}$ $0.060$ $0.060$
EncDec + DDPM $0.052$ $0.042$ $0.039$ $0.071$ $0.048$ $0.043$
EncDec + FM $0.051$ $\mathbf{0.041}$ $\mathbf{0.038}$ $0.070$ $\mathbf{0.047}$ $\mathbf{0.041}$
Table 2: Evaluations for different models on trajectory prediction with a 2 second horizon.
Different policy representations excel in different settings. In Table 2 we observe that the encoder-decoder flow matching (“FM”) model outperforms the other models for $K=5$ and $K=10$ by up to 34%. As expected, denoising diffusion (“DDPM”) and FM evaluations improve as $K$ increases, while behavior cloning (“BC”) remains the same independent of $K$ as it is deterministic. Note however that for the $K=1$ setting, BC outperforms both diffusion and flow-matching by about 15%. This suggests that the average prediction of BC is better than DDPM and FM, while the best prediction of DDPM and FM is better than BC’s single prediction.
Figure 4: Distance metrics w.r.t. training dataset size, where size is plotted on a log-scale. Performance improves as the dataset gets larger.
Performance degrades as the prediction horizon increases. In the remaining experiments we vary different properties while fixing the model to the simplest policy: decoder-only behavior cloning. In Table 3 we see that reducing the horizon from 2 seconds to 1 second improves average and final distance by 31% and 21% respectively, while increasing the horizon from 2 to 3 seconds worsens average and final distance by 18% and 11% respectively. Intuitively, accurate prediction becomes more challenging as the horizon increases as the model must predict 48-dimensional dexterous actions farther into the future.
Avg Distance (m) Final Distance (m)
Model $H=30$ (1s) $H=60$ (2s) $H=90$ (3s) $H=30$ (1s) $H=60$ (2s) $H=90$ (3s)
Dec + BC $\mathbf{0.031}$ $0.045$ $0.053$ $\mathbf{0.049}$ $0.062$ $0.069$
Table 3: Results for models trained and evaluated with different prediction horizons. As expected, accuracy falls as the prediction horizon increases. $H=60$ values are repeated from Table 2 for convenience.
Model Avg Distance (m) Final Distance (m)
Dec + BC $0.045$ $0.062$
Dec + BC w/ goal image $\mathbf{0.035}$ $\mathbf{0.029}$
Table 4: Visual goal-conditioning results. Training a model with visual goal conditioning reduces average distance by 22% and final distance by 53%.
Visual goal-conditioning significantly improves performance. In Table 4 we observe that visual goal-conditioning reduces average distance by 22% and final distance by 53%. Intuitively, a visual goal provides a visual “anchor” to ground the endpoint of the predicted trajectory and mitigate multimodality. This yields a baseline score for the inverse dynamics benchmark specified in Section 4.
Medium-size model capacity is sufficient for the current dataset size. We train and evaluate a larger Dec+BC model with 500 million parameters as opposed to the default 200 million parameters. The larger model attains average distance 0.045 and final distance 0.062, exactly the same as the default 200 million parameter model. This may increase accessibility for the EgoDex benchmarks, as medium-size models fit comfortably on commodity GPU hardware.
Performance scales with dataset size. In Figure 4 we observe that average and final distance improve as dataset size increases. Results suggest that performance scales with data, motivating the collection of large-scale egocentric datasets like EgoDex.
6 Research Use Cases
Robotics
Given a hypothetical robot embodiment that mimics the visual appearance, kinematics, and dynamics of human arms and hands, a control policy mapping egocentric video to 3D arm and hand pose deltas such as those trained in Section 5 may be deployable zero-shot without additional data or fine-tuning. Such an embodiment would lie within the training data distribution, both visually and dynamically.
While significant progress has been made in the development of robot hardware with humanoid morphologies and dexterous hands, there remains a prohibitive embodiment gap between humans and today’s robots. Some options for bridging the embodiment gap include 1) co-training with a small-scale robot dataset, as demonstrated by [15, 30]; 2) pretraining with large-scale human data and supervised fine-tuning with smaller-scale robot data, similar to the training recipe for large language models; 3) training a visual encoder on the human data for more data-efficient imitation learning downstream [22]; 4) learning robot manipulation priors from the human-object interaction trajectories and then fine-tuning with reinforcement learning or imitation learning [35, 10].
Perception
EgoDex can be used for learning tasks such as action recognition and human-object interaction detection from egocentric videos. Datasets like EPIC-KITCHENS [8] has demonstrated the value of egocentric video for recognizing and anticipating daily actions, and even more challenging tasks like detecting active objects and predicting state changes from egocentric video. Researchers can also study which objects are involved in each action and how. For example, one could model the contact points, grasps, and trajectories when using a tool (screwdriver, scissors, etc.). A related task is learning object affordances, i.e., understanding what actions each object supports.
Video Generation and World Models
Recent advances in large-scale diffusion models have significantly enhanced the capabilities of language-conditioned video generation, producing temporally consistent and semantically accurate visual narratives from natural language inputs [18, 29, 24]. These generative frameworks have demonstrated potential not only in creating realistic and detailed video content but also as world models for decision-making tasks, supporting reinforcement learning agents by simulating future outcomes based on predicted visual dynamics [40, 3, 41, 12]. Despite these impressive advancements, there remains a substantial research gap in video generation and world modeling from an egocentric viewpoint. Egocentric perspectives present unique challenges, including managing significant viewpoint variability, maintaining temporal and spatial consistency amid frequent camera movements, and accurately reflecting agent-centric interactions and intentions. Since EgoDex provides large-scale video data with 3D pose and language annotations, it enables the training of an egocentric world model.
7 Conclusion
We introduce EgoDex, a massive dataset of egocentric video paired with 3D pose annotations in a wide range of dexterous manipulation tasks. We train and evaluate imitation learning policies for hand trajectory prediction on this data.
While EgoDex has significant diversity across tasks and manipulation behaviors, it is limited in background and scene diversity. The dexterous annotations can also be imperfect, especially during heavy occlusion (e.g., towel folding) or very high speed motions, as they are themselves model predictions. Future work involves procedural background randomization on the existing data [42] as well as data collection in more diverse environments.
Appendix A Appendix
A.1 Complete List of Tasks
We provide a complete list of task names here, labeled as they appear in the dataset and separated by task type (reversible, reset-free, or reset, with definitions in Section 3.4). Recall that each reversible task is actually a pair of two tasks. There are a total of 194 tasks. See Figure 5 for a visual of a subset of the objects used in the various manipulation tasks.
For users interested in robot deployment, the basic_pick_place task in particular has a large amount of very diverse pick-and-place data as well as high-quality language annotation.
Reversible (76 $\times$ 2 total tasks):
  • braid_unbraid
  • charge_uncharge_airpods
  • deal_gather_cards
  • fry_bread
  • assemble_disassemble_furniture_bench_chair
  • assemble_disassemble_furniture_bench_drawer
  • assemble_disassemble_furniture_bench_square_table
  • fold_unfold_paper_basic
  • insert_remove_furniture_bench_cabinet
  • gather_roll_dice
  • insert_remove_airpods
  • insert_remove_drawer
  • insert_remove_shirt_in_tube
  • insert_remove_usb
  • load_dispense_ice
  • open_close_insert_remove_tupperware
  • pick_up_and_put_down_case_or_bag
  • put_away_set_up_board_game
  • screw_unscrew_fingers_fixture
  • sleeve_unsleeve_cards
  • stack_unstack_cups
  • thread_unthread_bead_necklace
  • tie_and_untie_shoelace
  • insert_remove_tennis_ball
  • open_close_insert_remove_case
  • pick_place_food
  • put_in_take_out_glasses
  • screw_unscrew_allen_fixture
  • set_up_clean_up_chessboard
  • slot_batteries
  • stack_unstack_bowls
  • stack_unstack_tupperware
  • throw_collect_objects
  • vertical_pick_place
  • wash_put_away_dishes
  • add_remove_lid
  • arrange_topple_dominoes
  • assemble_disassemble_legos
  • assemble_disassemble_soft_legos
  • assemble_disassemble_structures
  • assemble_disassemble_tiles
  • boil_serve_egg
  • build_unstack_lego
  • charge_uncharge_device
  • clip_unclip_papers
  • crumple_flatten_paper
  • fry_egg
  • assemble_disassemble_furniture_bench_desk
  • assemble_disassemble_furniture_bench_lamp
  • assemble_disassemble_furniture_bench_stool
  • fold_stack_unstack_unfold_cloths
  • fold_unfold_paper_origami
  • insert_remove_furniture_bench_round_table
  • insert_remove_bagging
  • insert_remove_cups_from_rack
  • insert_remove_plug_socket
  • insert_remove_utensils
  • lock_unlock_key
  • open_close_insert_remove_box
  • scoop_dump_ice
  • screw_unscrew_bottle_cap
  • setup_cleanup_table
  • stock_unstock_fridge
  • stack_unstack_plates
  • throw_and_catch_ball
  • tie_untie_rubberband
  • wrap_unwrap_food
  • zip_unzip_bag
  • zip_unzip_case
  • assemble_disassemble_jigsaw_puzzle
  • stack_unstack_tetra_board
  • stack_remove_jenga
  • insert_dump_blocks
  • rake_smooth_zen_garden
  • play_reset_connect_four
  • insert_remove_bookshelf
Reset-free (28 total tasks):
  • color
  • fidget_magnetic_spinner_rings
  • measure_objects
  • staple_paper
  • use_rubiks_cube
  • wash_kitchen_dishes
  • wipe_screen
  • knead_slime
  • point_and_click_remote
  • type_keyboard
  • clean_surface
  • dry_hands
  • play_mancala
  • flip_coin
  • flip_pages
  • paint_clean_brush
  • play_piano
  • push_pop_toy
  • put_toothpaste_on_toothbrush
  • wash_fruit
  • wipe_kitchen_surfaces
  • stamp_paper
  • blowdry_hair
  • knit_scarf
  • makeup
  • write
  • clean_cups
  • roll_ball
Reset (14 total tasks):
  • clean_tableware
  • declutter_desk
  • basic_pick_place
  • stack
  • make_sandwich
  • peel_place_sticker
  • sweep_dustpan
  • wrap
  • assemble_jenga
  • basic_fold
  • pour
  • sort_beads
  • use_chopsticks
  • play_reversi
Figure 5: Some of the objects used in the various manipulation tasks.
A.2 Complete List of Skeletal Joints
The annotations consist of SE(3) poses (represented as 4 $\times$ 4 homogeneous transformation matrices) for each of the following joints, labeled by their names as they appear in the dataset:
Upper Body:
hip, spine1, spine2, spine3, spine4, spine5, spine6, spine7, neck1, neck2, neck3, neck4, leftShoulder, leftArm, leftForearm, leftHand, rightShoulder, rightArm, rightForearm, rightHand
Left Hand:
leftIndexFingerIntermediateBase, leftIndexFingerIntermediateTip, leftIndexFingerKnuckle, leftIndexFingerMetacarpal, leftIndexFingerTip, leftLittleFingerIntermediateBase, leftLittleFingerIntermediateTip, leftLittleFingerKnuckle, leftLittleFingerMetacarpal, leftLittleFingerTip, leftMiddleFingerIntermediateBase, leftMiddleFingerIntermediateTip, leftMiddleFingerKnuckle, leftMiddleFingerMetacarpal, leftMiddleFingerTip, leftRingFingerIntermediateBase, leftRingFingerIntermediateTip, leftRingFingerKnuckle, leftRingFingerMetacarpal, leftRingFingerTip,
leftThumbIntermediateBase, leftThumbIntermediateTip, leftThumbKnuckle, leftThumbTip
Right Hand:
rightIndexFingerIntermediateBase, rightIndexFingerIntermediateTip, rightIndexFingerKnuckle, rightIndexFingerMetacarpal, rightIndexFingerTip, rightLittleFingerIntermediateBase, rightLittleFingerIntermediateTip, rightLittleFingerKnuckle, rightLittleFingerMetacarpal, rightLittleFingerTip, rightMiddleFingerIntermediateBase,
rightMiddleFingerIntermediateTip, rightMiddleFingerKnuckle, rightMiddleFingerMetacarpal, rightMiddleFingerTip, rightRingFingerIntermediateBase, rightRingFingerIntermediateTip, rightRingFingerKnuckle, rightRingFingerMetacarpal, rightRingFingerTip, rightThumbIntermediateBase, rightThumbIntermediateTip, rightThumbKnuckle, rightThumbTip
Note that leftHand and rightHand refer to the wrists. Note also that the joint confidence values in the data behave differently for the wrists and the hands. Wrist confidence values (for leftHand and rightHand) indicate whether each hand is detected as a whole, while finger joint confidence values indicate confidence relative to the wrist. If, for instance, the left index fingertip has high confidence but the left wrist has low confidence, it is unlikely that the left index fingertip is reliable.
A.3 Training Details
In the experiments section we train and evaluate 14 different models: 6 combinations of architectures and policy optimization methods, 4 additional models with different training dataset sizes, 2 additional models with different prediction horizons, 1 model with a larger model size, and 1 model with visual goal-conditioning. See Figure 6 for intuition on the model architecture.
Each model is trained and evaluated on a single node with 96 logical CPUs (48 physical CPUs) and 8 NVIDIA A100 GPUs each with 80GB RAM. Training is run for 50,000 gradient steps with a batch size of 2048 (256 per GPU with data parallelism), at which point training and validation loss plateau. The full training run takes approximately 72 hours. The models are optimized with Adam and a learning rate of 1e-4. Image observations are resized to 224 $\times$ 224 and sent through a pretrained ResNet encoder, while language annotations are passed through a frozen CLIP [31] encoder. Only the current image observation and proprioceptive state are passed as input to the policy (i.e., no history); adding history may improve performance. DDPM and FM models are trained and evaluated with 16 sampling steps. All other hyperparameters are the defaults from the X-IL codebase [14].
Figure 6: Model architectures.
A.4 Dataset Access
The full dataset as well as sample Python code for data loading are available at https://github.com/apple/ml-egodex. Instructions for downloading the dataset are in the README.md, and repeated below for convenience.
The dataset is split into a training set, test set, and additional data collected after the initial 99/1 train/test split was frozen for all experimental results. These have about 725 hours, 7 hours, and 97 hours, respectively. The training set is further divided into 5 zip files for portability, each with a size of around 300 GB. The data is publicly available for download at the following URL: https://ml-site.cdn-apple.com/datasets/egodex/[filename].zip, where [filename] is one of part1, part2, part3, part4, part5, test, extra (for example, https://ml-site.cdn-apple.com/datasets/egodex/test.zip). The 5 parts belong to the training set, whereas test.zip is the test set and extra.zip is the additional data. There is also a README.md under the same prefix with further instructions on the dataset structure. The dataset is licensed under CC-by-NC-ND terms.